# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

### v1

@@ II-C Related Topics
Evolutionary Robotics (ER) originally focused on evolving robot behaviors [bongard2013evolutionary, doncieux2015evolutionary, medvet2023evolutionary]. For example, behavioral trees have been evolved using genetic programming [scheper2016behavior]. Over time, the scope expanded to include the evolutionary design of both robotic hardware and control policies. A seminal contribution is Sims’ work [sims1994evolving], which evolves 3D virtual creatures by jointly optimizing their graph-based morphologies and neural controllers. Contemporary ECD builds on these insights but extends beyond traditional evolutionary robotics, incorporating a wider range of optimization and learning techniques to improve scalability, generalization, and applicability.
Artificial Life (A-Life) investigates the co-evolution of morphology and control, aiming to recreate life-like processes in virtual environments [moreno2005agency, aguilar2014past]. While A-Life shares the central concept of brain–body–environment coupling with ECD, it prioritizes the exploration of biological phenomena such as self-organization [gershenson2020self], adaptation [bedau2003artificial], and evolutionary dynamics [ito2016population, pilat2012evolution] in nature, rather than engineering practical embodied agents with specific functionalities.
Computational Robot Design leverages algorithmic tools to automate the synthesis of robots, typically following a sequential design process. By formulating complex constrained optimization problems, this approach employs techniques such as topology optimization [sigmund2000topology, sigmund2013topology] and CAD integration [tian2022assemble, willis2022joinable] to address structural feasibility, material selection, electronics integration, fabrication constraints, and dynamic modeling. Control, however, is often simplified or treated as a secondary component. In contrast, ECD is a broader concept that tightly couples morphology and control within a unified optimization loop, aiming not only to generate feasible body designs but also to produce adaptive, task-capable behaviors.
Robot Learning aims to equip robots with the ability to autonomously acquire perceptual, decision-making, and action skills, rather than relying solely on hand-engineered solutions [kroemer2021review]. It is closely connected to ECD, as learning-based control optimization serves as a central mechanism for co-design in tasks such as locomotion, manipulation, and navigation. Recent advances in robot learning have integrated model-based approaches [grandia2023perceptive], reinforcement learning [matsuo2022deep, singh2022reinforcement], large-model-based reasoning [xiao2025robot], imitation learning [hussein2017imitation, fang2019survey], and self-supervised methods [nair2020contextual, jeong2020self, kahn2021badgr] to enhance high-dimensional sensorimotor loops. However, most robot learning frameworks assume a fixed morphology. ECD generalizes this setting by treating morphology and control as interdependent variables and optimizing them jointly.
@@ III Taxonomy: Decomposing Embodied Co-Design
Figure 3: An overview of the embodied co-design landscape including foundational elements and co-design frameworks.
ECD operates within an expansive workspace shaped by the interplay of control, morphology, and task environment. As shown in Fig. 3, the lower layer of our taxonomy organizes this space by complexity, while the upper layer defines four ECD frameworks, each with sub-categories that guide the co-design process. In the following sections, we detail each component, highlight recent advances, discuss common constraints, summarize key evaluation metrics, and illustrate a representative ECD pipeline using a prototype.
@@ III-A1 Design Space
The design space encompasses both the ways controllers are parameterized and the modes in which they operate.
Controller Parameters can be predefined—such as action primitives [iii2021taskagnostic], fixed waveform generators [cheney2013unshackling], and hand-crafted gaits [kobayashi2024computational]—or be fully learnable, optimized through methods such as trajectory optimization [spielberg2017functional, bergonti2024co], reinforcement learning, and imitation learning. Modern ECD approaches typically employ learnable controllers, which enable adaptation to varying morphologies and support flexible, real-time improvements in behavior.
@@ III-A2 Representation
The controller representation defines the mapping between the control policy and the agent’s physical structure. Classical approaches adopt either individual or universal representations, with recent research exploring hybrid variants that combine their strengths.
@@
Individual representations directly bind the controller architecture or dimensionality to a specific morphology. Examples include PID controllers with per-joint gains that account for the inertia and stiffness of specific body segments [ang2005pid, borase2021review], as well as Model Predictive Control (MPC) when an accurate dynamic model is available [schwenzer2021review, carron2019data]. Another example is Central Pattern Generators (CPGs), which mirror a robot’s limb topology to generate coordinated rhythmic motion [ijspeert2008central, bliss2012central, deshpande2023deepcpg]. Higher-level neural controllers, such as Multi-Layer Perceptrons (MLPs), adopt input/output sizes tied to a robot’s sensors and actuators.
Individualized controllers are appropriate for topology-fixed agents that share the same morphological structures but differ solely in their kinematic or dynamic parameters. This specialization, however, reduces transferability: even minor morphological changes—such as the addition of a joint or modifications to limb geometry—typically require controller restructuring or retraining. Consequently, per-agent representations struggle to generalize across heterogeneous state and action spaces, restricting scalability.
@@
Universal representations aim to generalize across diverse morphologies. Representing a body as a kinematic graph naturally invites Graph Neural Networks (GNNs) [munikoti2023challenges]: their message-passing mechanisms enable a single policy to generalize across graphs with varying node counts, thereby supporting scalable control across different robotic embodiments. Representative GNN-based controllers include Nervernet [wang2018nervenet], DGN [pathak2019learning], Whitman et al. [whitman2023learning], SMP [huang2020one], MAT [li2024mat], MG [pan2025morphology], and GCNT [ijcai2025p972]. Subsequent research in cognitive science has further shown that modularity can enhance control generalization [schilling2024modularity, nadizar2022collective, pigozzi2022evolving, pigozzi2022evolving2, dong2022low, zhang2024composer, tjandrasuwita2024memo]. To capture the long-range dependencies that GNNs inherently struggle with, transformer-based controllers incorporate global attention over all body parts. This idea has been explored in numerous embodied AI systems, such as Amorpheus [kurin2020my], MetaMorph [gupta2022metamorph], SWAT [hong2021structure], AnyMorph [trabucco2022anymorph], Xiong et al. [xiong2023universal], Body Transformer [sferrazza2024body], HeteroMorpheus [hao2024heteromorpheus], and URMA [bohlinger2024one]. Beyond ECD, universal controllers have also shown strong utility in transfer learning [chen2018hardware, peng2018sim, cho2024meta], meta-learning [nagabandi2018learning, ghadirzadeh2021bayesian], and cross-embodiment learning [yang2024pushing, shafiee2024manyquadrupeds, wang2024sparse, chen2024mirage, patel2025get, mertan2024towards, feng2023genloco, luo2024moral]. This underscores their potential as a unifying representation for embodied AI.
@@
Hybrid representations seek to balance specialization and generality by decomposing the controller or control mechanism into a shared, morphology-agnostic module that encodes universal motor and perceptual priors [bjelonic2023learning], and a morphology-specific module that adapts outputs to specific embodiments [chen2023evolving, luck2020data, luck2021robot, sliacka2023co, sun2023co]. Hybrid representations can be seamlessly integrated into various pipelines, such as pretraining-finetuning [yue2025toward, jin2025evolutionary] and meta-learning [belmonte2022meta], to facilitate cross-morphology generalization and faster adaptation during the ECD process.
@@ III-A3 Discussion
Training a dedicated controller for each morphology yields an accurate estimate of that body’s potential by fully exploiting its passive dynamics and affordances. However, this approach is computationally expensive and biases the search toward morphologies that are easier to learn. Empirical results from [kvalsund2022centralized, mertan2023modular, chiappa2022dmap, luo2023comparison] have demonstrated that universal controllers can promote better synchronization, greater morphological diversity, and reduced premature convergence. However, they also require more comprehensive evaluation metrics to “diagnose” whether performance gains arise from enhanced controller capabilities or advantageous morphological adaptations.
These observations raise an important question: must one optimize a controller to evaluate morphology? Recent work in ECD suggests that the answer may be no. For instance, TAME [iii2021taskagnostic] introduces an information-theoretic, controller-free fitness function that measures a morphology’s intrinsic ability to explore its environment and produce predictable state transitions using only random action primitives.
@@ III-B1 Design Space
Morphological Parameters are commonly categorized as discrete, continuous, and hybrid. Discrete parameters typically arise in modular systems where agents are composed of a finite set of building blocks, topologies, or design configurations. Continuous parameters govern real-valued attributes such as link lengths, shape, and material properties. More general embodied systems adopt a hybrid scheme in which certain choices, such as whether to attach a limb, are discrete, whereas attributes like limb length are continuous. This formulation balances structural flexibility with optimization tractability, enabling scalable search over high-dimensional morphologies.
Topological Structure further defines the scope of morphological exploration. In fixed topology settings, such as those involving bipedal [ha2019reinforcement], quadrupedal [belmonte2022meta], or manipulator skeletons [cai2023task2morph], optimization is constrained to parametric variations within a predefined template. This approach ensures stability and facilitates hardware deployment, but it restricts the discovery of novel morphologies. In contrast, unfixed topology settings allow for arbitrary structural modifications, ranging from voxel-based [bhatia2021evolution] to graph-based agents [yuan2022transformact] capable of growing, pruning, or reconfiguring their skeletons. While enabling open-ended exploration and emergent specialization, these approaches also raise concerns regarding physical plausibility, degeneracy, and manufacturability.
@@ III-B2 Representation
Indirect encodings compress the agent’s morphological information into compact descriptors or generative rules, which are then expanded into complete morphologies through developmental or decision-making processes. Examples include L-Systems [veenstra2022effects], generative grammars [zhao2020robogrammar, zhao2022automatic, zharkov2024synergizing], Neural Cellular Automata (NCA) [veenstra2020different, wang2023curriculum], Compositional Pattern-Producing Networks (CPPNs) [cheney2013unshackling], and reinforcement learning-based decision-making [yuan2022transformact, wang2023curriculum]. Indirect encodings typically yield smoother fitness landscapes [thomson2024understanding], facilitate the escape from morphological local optima, and produce modular, symmetric morphologies with enhanced transferability.
Hybrid encodings aim to combine the strengths of both approaches by explicitly parameterizing part of the morphology while generating the remainder procedurally or via learned models [dongleveraging2024, ringel2025text2robot, saito2024effective, le2024improving, wang2023softzoo, mertan2025controller]. This design facilitates both global exploration and local refinement. However, it also introduces the challenge of balancing the search across heterogeneous spaces while maintaining coherence between explicit and generated components.
@@ III-B3 Discussion
The trade-offs between direct and indirect morphological encodings remain an open question in ECD. Prior work [pigozzi2023morphology] has reported that direct encodings often yield faster learning and more consistent convergence. In contrast, indirect encodings tend to generate more diverse, spatially coherent, and globally varying morphologies, although the corresponding learning improvements can be less predictable. Comparative studies [komosinski2001comparison, veenstra2017evolution] further demonstrated that indirect encodings are advantageous during early exploration due to their compactness and genetic reuse, whereas direct encodings may outperform them in later optimization stages. These insights underscore the potential of employing hybrid morphology representation strategies that begin with indirect encodings for broad exploration, followed by a transition to direct encodings for fine-grained refinement.
@@ III-C1 Design Space
The design space of the task environment specifies whether environmental parameters remain fixed or vary during co-design, and how feedback signals are delivered.
Environmental Parameters are typically categorized into predefined and modifiable types. In predefined settings, all environmental attributes, such as terrain roughness and object placement, are fixed, resulting in stable and repeatable dynamics. Such static environments are widely utilized to isolate environmental factors, thereby enabling the targeted specialization of morphology and control. However, this specialization often results in brittle solutions that degrade under unmodeled perturbations or deployment beyond training conditions. In contrast, treating the environment as modifiable—for example, by sampling from a distribution of scenario parameters, introducing curriculum-based difficulty [wang2019poet], or employing non-stationary task sequences—encourages the emergence of agents that generalize more effectively [bossens2020qed, auerbach2014environmental, urbina2024brain].
Observability is a fundamental property of the task environment, as it determines the fidelity and completeness of the information an agent can perceive about the underlying system state. In a fully observable setting, the agent’s observations constitute a sufficient statistic of the true environment state. This condition simplifies learning and planning, as seen in classical control and various reinforcement learning tasks. Conversely, partial observability arises when agents are provided with incomplete [schaff2022soft], noisy [bjelonic2023learning], or local per-body/per-joint information [wang2023curriculum, schaff2022n, yuan2022transformact], which decouples immediate observations from the global state. Under such conditions, the agent must maintain an internal belief state or memory to integrate temporal observations. The degree of observability, therefore, shapes an agent’s perceptual and control mechanisms, ranging from simple reactive policies to sophisticated memory architectures or active perception strategies. It is also a critical axis for benchmarking an agent’s robustness and generalization in real-world scenarios.
@@ III-C2 Representation
This defines the objectives, constraints, and evaluative (reward) signals that guide both the learning dynamics and the emergent morphology–control solutions in ECD. We broadly categorize these representations into three types: task-driven, task-agnostic, and hybrid.
@@ III-C3 Discussion
The task environment shapes, and is shaped by, the agent’s morphology and control. Prior studies have demonstrated that robots co-designed on uneven terrain tend to develop more stable, balanced, and limb-rich morphologies, whereas those created on flat terrain favor simpler forms [miras2020environmental, miras2019effects, spanellis2021environment, corucci2018evolving, ito2013alife]. Similarly, morphologies that evolved in complex, multi-objective environments transfer more effectively to novel tasks than those from simpler settings [gupta2021embodied]. Moreover, multi-task scenarios in ECD introduce additional brain–body trade-offs, where improvements on one task may degrade performance on others [wang2023preco]. As a result, multi-task co-design aims to cultivate broadly capable brain–body pairs by leveraging shared structures across tasks rather than optimizing each task independently.

### v2

@@ II-C Related Topics
Evolutionary Robotics (ER) originally focused on the evolution of robot behaviors [bongard2013evolutionary, scheper2016behavior]. Over time, the scope expanded to include the evolutionary design of both robotic hardware and control policies. A seminal contribution is Sims’ work [sims1994evolving], which evolves 3D virtual creatures by jointly optimizing their graph-based morphologies and neural controllers. Contemporary ECD builds upon these insights while extending beyond traditional ER by incorporating a broader array of representation and learning techniques to enhance scalability, generalization, and applicability.
Artificial Life (A-Life) investigates the co-evolution of morphology and control, aiming to recreate life-like processes in virtual environments [moreno2005agency, aguilar2014past]. While A-Life shares the central concept of brain–body–environment coupling with ECD, it prioritizes the exploration of biological phenomena such as self-organization [gershenson2020self], adaptation [bedau2003artificial], and evolutionary dynamics [ito2016population] in nature, rather than engineering practical embodied agents with specific functionalities.
Computational Robot Design leverages algorithmic tools to automate the synthesis of robots, typically following a sequential design process. This approach formulates complex constrained optimization problems and employs techniques such as topology optimization [sigmund2000topology, sigmund2013topology] and CAD integration [tian2022assemble, willis2022joinable] to address issues related to structural feasibility, material selection, electronics integration, and fabrication constraints. Control, however, is often simplified or regarded as a secondary component. In contrast, ECD is a broader concept that tightly integrates morphology and control within a unified optimization loop, aiming to generate not only feasible body designs but also adaptive, task-capable behaviors.
Robot Learning aims to equip robots with the ability to autonomously acquire perceptual, decision-making, and action skills, rather than relying solely on hand-engineered solutions [kroemer2021review]. It is closely connected to ECD, as learning-based control optimization serves as a central mechanism for solving tasks such as locomotion, manipulation, and navigation. Recent advances in robot learning have integrated model-based approaches [grandia2023perceptive], reinforcement learning [matsuo2022deep], large-model-based reasoning [xiao2025robot], imitation learning [hussein2017imitation], and self-supervised methods [jeong2020self] to enhance high-dimensional sensorimotor loops. However, most robot learning frameworks assume a fixed morphology. ECD generalizes this setting by treating morphology and control as interdependent variables and optimizing them jointly.
@@ III Taxonomy: Decomposing Embodied Co-Design
Fig. 2: An overview of the embodied co-design landscape including foundational elements and co-design frameworks.
ECD operates within an expansive workspace shaped by the interplay of control, morphology, and task environment. As shown in Fig. 2, our taxonomy organizes these fundamental elements into a lower layer, while the upper layer defines four ECD frameworks, each with sub-categories that guide the co-design process. In the following sections, we detail each component, highlight recent advances, discuss common constraints, summarize key evaluation metrics, then illustrate a representative ECD pipeline using a prototype.
@@ III-A1 Design Space
The design space of controlling brain encompasses both the ways by which controllers are parameterized and the modes in which they operate.
Controller Parameters can be predefined—such as action primitives [iii2021taskagnostic], fixed waveform generators [cheney2013unshackling], and hand-crafted gaits [kobayashi2024computational]—or they can be fully learnable, optimized through methods such as trajectory optimization, reinforcement learning, and imitation learning. Modern ECD approaches typically employ learnable controllers that enable adaptation to varying morphologies and support flexible, real-time improvements in behavior.
@@ III-A2 Representation
The controller representation defines the mapping between the control policy and the agent’s physical structure. Classical approaches adopt either individual or universal representations, while recent research explores hybrid variants that combine their strengths.
@@
Individual representations directly bind the controller architecture or dimensionality to a specific morphology. Examples include low-level PID controllers with per-joint gains that account for the inertia and stiffness of specific body segments [ang2005pid, borase2021review], as well as Model Predictive Control (MPC) when an accurate dynamic model is available [schwenzer2021review, carron2019data]. Another example is Central Pattern Generators (CPGs), which mirror a robot’s limb topology to produce coordinated rhythmic motion [bliss2012central, deshpande2023deepcpg]. Higher-level neural controllers, such as Multi-Layer Perceptrons (MLPs), adopt input/output sizes tied to a robot’s sensors and actuators. Individualized controllers are suitable for topology-fixed agents that share identical morphological structures but differ solely in their kinematic or dynamic parameters. This specialization, however, reduces transferability; even minor morphological changes—such as the addition of a joint or modifications to limb geometry—typically necessitate restructuring the controller. Consequently, per-agent representations struggle to generalize across heterogeneous state and action spaces.
@@
Universal representations aim to generalize across diverse morphologies. Representing a body as a kinematic graph naturally invites the use of Graph Neural Networks (GNNs) [munikoti2023challenges]; their message-passing mechanisms enable a single policy to generalize across graphs with varying node counts, thereby supporting scalable control across different robotic embodiments. Representative GNN-based controllers include Nervernet [wang2018nervenet], DGN [pathak2019learning], Whitman et al. [whitman2023learning], SMP [huang2020one], MAT [li2024mat], MG [pan2025morphology], and GCNT [ijcai2025p972]. Subsequent research in cognitive science has further demonstrated that modularity can enhance control generalization [schilling2024modularity]. To capture the long-range dependencies that GNNs inherently struggle to address, transformer-based controllers incorporate global attention over all body parts. This concept has been explored in numerous embodied AI studies, such as Amorpheus [kurin2020my], MetaMorph [gupta2022metamorph], SWAT [hong2021structure], AnyMorph [trabucco2022anymorph], Xiong et al. [xiong2023universal], Body Transformer [sferrazza2024body], HeteroMorpheus [hao2024heteromorpheus], and URMA [bohlinger2024one]. Beyond ECD, universal controllers have also shown strong utility in transfer learning [chen2018hardware, peng2018sim, cho2024meta], meta-learning [nagabandi2018learning, ghadirzadeh2021bayesian], and cross-embodiment learning [yang2024pushing, shafiee2024manyquadrupeds, wang2024sparse, chen2024mirage, patel2025get, mertan2024towards, feng2023genloco, luo2024moral].
@@
Hybrid representation seeks to balance specialization and generalization by decomposing the controller or control mechanism into a shared, morphology-agnostic module that encodes universal motor and perceptual priors, and a morphology-specific module that adapts outputs to specific embodiments [bjelonic2023learning, chen2023evolving, luck2020data, luck2021robot, sliacka2023co, sun2023co]. Hybrid representations can be seamlessly integrated into various pipelines, such as pretraining-finetuning [yue2025toward, jin2025evolutionary] and meta-learning [belmonte2022meta], to facilitate cross-morphology generalization and accelerate adaptation during the ECD process.
@@ III-A3 Discussion
Training a dedicated controller can provide an accurate estimate of a morphology’s potential by fully exploiting its passive dynamics and affordances. Empirical results from [kvalsund2022centralized, mertan2023modular, chiappa2022dmap, luo2023comparison] have demonstrated that well-designed controllers can enhance morphological diversity and mitigate premature convergence. However, they require more comprehensive evaluation efforts to “diagnose” whether performance gains arise from enhanced controller capabilities or advantageous morphological adaptations. This raises an important question: must one optimize a controller to evaluate morphology? Recent work in ECD suggests that the answer may be no. For instance, TAME [iii2021taskagnostic] introduces an information-theoretic, controller-free fitness function that measures a morphology’s intrinsic ability to explore its environment and produce predictable state transitions using only random action primitives.
@@ III-B1 Design Space
Morphological Parameters are commonly categorized as discrete, continuous, and hybrid. Discrete parameters typically arise in modular systems where agents are composed of a finite set of building blocks, topologies, or design configurations. Continuous parameters govern real-valued attributes such as link lengths, shapes, and material properties. More general embodied systems adopt a hybrid scheme in which certain choices, such as whether to attach a limb, are discrete, whereas attributes like limb length are continuous. This formulation balances structural flexibility with optimization tractability, enabling scalable searches over high-dimensional morphologies.
Topological Structure further defines the scope of morphological exploration. In fixed topology settings, such as those involving bipedal [ha2019reinforcement], quadrupedal [belmonte2022meta], or manipulator skeletons [cai2023task2morph], optimization is constrained to parametric variations within a predefined template. This approach ensures stability and facilitates hardware deployment; however, it restricts the discovery of novel morphologies. In contrast, unfixed topology settings permit arbitrary structural modifications, ranging from voxel-based [bhatia2021evolution] to graph-based agents [yuan2022transformact] capable of growing, pruning, or reconfiguring their skeletons. While enabling open-ended exploration, these approaches also raise concerns regarding manufacturability.
@@ III-B2 Representation
Indirect encodings compress the agent’s morphological information into compact descriptors or generative rules, which are then expanded into complete morphologies through developmental or decision-making processes. Examples include L-systems [veenstra2022effects], generative grammars [zhao2020robogrammar], Neural Cellular Automata (NCA) [veenstra2020different], Compositional Pattern-Producing Networks (CPPNs) [cheney2013unshackling], and reinforcement learning-based decision-making [yuan2022transformact, wang2023curriculum]. Indirect encodings typically yield smoother fitness landscapes [thomson2024understanding], facilitate the escape from morphological local optima, and produce modular, symmetric morphologies with enhanced transferability.
Hybrid encodings aim to combine the strengths of both approaches by explicitly parameterizing part of the morphology while generating the remainder procedurally or via learned models [dongleveraging2024, ringel2025text2robot, saito2024effective, le2024improving, mertan2025controller]. This design facilitates both global exploration and local refinement. However, it also introduces the challenge of balancing the search across heterogeneous spaces while maintaining coherence between explicit and generated components.
@@ III-B3 Discussion
The trade-off between direct and indirect morphological encodings remains an open question in ECD. Prior work [pigozzi2023morphology] has reported that direct encodings often yield faster learning and more consistent convergence. In contrast, indirect encodings generate more diverse, spatially coherent, and globally varying morphologies; however, their impact on learning performance is less predictable. Comparative studies [komosinski2001comparison, veenstra2017evolution] further demonstrate that indirect encodings provide advantages during early exploration due to their compactness and genetic reuse, while direct encodings may excel during later optimization stages. These insights underscore the potential of hybrid strategies that begin with indirect encodings for broad exploration and subsequently transition to direct encodings for fine-grained refinement.
@@ III-C1 Design Space
The design space of task environment specifies whether environmental parameters remain fixed or vary during co-design and how feedback signals are delivered.
Environmental Parameters are typically categorized into predefined and modifiable types. In predefined settings, all environmental attributes, such as terrain roughness and object placement, are fixed, resulting in stable and repeatable dynamics. However, this specialization often results in brittle solutions that degrade under unmodeled perturbations or when deployed beyond training conditions. In contrast, treating the environment as modifiable—for example, by sampling from a distribution of scenario parameters, introducing curriculum-based difficulty, or employing non-stationary task sequences—encourages agents to generalize more effectively.
Observability is a fundamental property of the task environment, as it determines the fidelity and completeness of the information an agent can perceive about the underlying system state. In a fully observable setting, the agent’s observations constitute a sufficient statistic of the true environment state. This condition simplifies learning and planning, as observed in classical control and various reinforcement learning tasks. Conversely, partial observability occurs when agents receive incomplete [schaff2022soft], noisy [bjelonic2023learning], or local per-body/per-joint information [wang2023curriculum, schaff2022n, yuan2022transformact], which decouples immediate observations from the global state. In such cases, the agent must maintain an internal belief state or memory to integrate temporal observations. The degree of observability, therefore, inherently influences an agent’s perceptual and control mechanisms; it is also a critical axis for benchmarking an agent’s robustness and generalization in real-world scenarios.
@@ III-C2 Representation
The environmental representation defines the objectives, constraints, and evaluative (reward) signals that guide the optimization dynamics of ECD. We broadly categorize these representations into three types: task-driven, task-agnostic, and hybrid.
@@ III-C3 Discussion
The task environment shapes, and is shaped by, the agent’s morphology and control. Prior studies indicate that robots co-designed on uneven terrain tend to develop more stable, balanced, and limb-rich morphologies, whereas flat terrains favor simpler forms [miras2020environmental, miras2019effects, spanellis2021environment, corucci2018evolving]. Moreover, morphologies designed in complex, multi-task environments exhibit stronger transfer to novel tasks than those developed in simpler settings [gupta2021embodied, wang2023preco]. Note that multi-task scenarios introduce additional brain–body trade-offs, as morphological adaptations that improve one task may degrade performance on others. Consequently, multi-task ECD seeks to cultivate broadly capable brain–body pairs by leveraging the shared structures across tasks.